# Saliency strikes back: How filtering out high frequencies improves white-box explanations

### v3

@@ 5 FORGrad: FOurier Reparation of the Gradients
With FORGrad, we propose to filter out high-frequency content, considered to be noise, to improve the XAI performance of white-box attribution methods. We therefore apply a low-pass filter on the Fourier spectrum of the gradient of the classifier w.r.t the input (i.e. $\nabla_{\bm{x}}\bm{f}(\bm{x})$). This low-pass filter is equipped with a mechanism to adjust the cut-off frequency, $\sigma^{\star}$, based on the specific network architecture and attribution method.
@@
Fig. 5B illustrates the sensitivity of the faithfulness to changes in the cut-off frequency, $\sigma$, underscoring the necessity of selecting an optimal $\sigma^{\star}$. In Fig. 5A, we provide visual examples using the Gradient Input method (Ancona et al., 2018), where the gradient is low-pass filtered at various cut-off frequencies. One can observe that the initial noisy patterns (Fig. 5A, $\sigma=220$ ) progressively evolve into clearer, and larger patches (Fig. 5A, $\sigma=80$), culminating in saliency maps that effectively emphasize the critical features for object recognition (Fig. 5A, $\sigma=10$). However, at very low $\sigma$ values (Fig. 5A, $\sigma=5$), the attribution map becomes overly diffuse, reducing interpretability. Fig. 5D compares the optimal $\sigma^{\star}$ values across different ResNetV2 (He et al., 2016b) architectures and attribution methods. We observe that even with the same attribution method, distinct architectures achieve their maximum faithfulness at different $\sigma^{\star}$ values. Likewise, the same network architecture can yield different optimal $\sigma^{\star}$ values when leveraging different attribution methods. This variability highlights the crucial role of a specific cut-off frequency, emphasizing that a one-size-fits-all approach may not be effective in maximizing the faithfulness of attribution methods in different contexts. Additionally, Fig. 5C presents a side-by-side comparison of attribution maps generated using the SmoothGrad method (Smilkov et al., 2017) (top-row) versus those refined with the FORGrad method (bottom row). A qualitative analysis shows that the FORGrad method yields more interpretable explanations, with a sharper focus on crucial features for object recognition. We refer the reader to Fig. A.10 for more examples on all white-box methods.
@@
Faith.($\uparrow$) $\mu$Fid.($\uparrow$) Time($\downarrow$)
White-box Saliency(Simonyan et al., 2013) 0.18 0.40 0.78
Saliency 0.27 0.39 0.89
Guidedbackprop(Ancona et al., 2018) 0.31 0.45 8.25
@@
Black & Gray-box GradCAM(Selvaraju et al., 2017a) 0.31 0.40 5.24
GradCAM++(Chattopadhay et al., 2018) 0.33 0.43 4.61
Occlusion(Ancona et al., 2018) 0.17 0.39 368
HSIC(Novello et al., 2022) 0.33 0.47 456
Sobol(Fel et al., 2021b) 0.34 0.47 578
RISE(Petsiuk et al., 2018) 0.41 0.34 626
@@
FORGrad allows white-box methods to rival with black-box methods: We now evaluate the FORGrad method using XAI benchmarks, focusing on two established metrics: Faithfulness and $\mu$Fidelity (detailed in section  3). Our findings, presented in Table 3, compare the original performance of white-box attribution methods, their enhanced versions via the FORGrad method (marked with a $\star$), black-box, and gray-box methods. This comparison also includes the computational time, measured on 100 images with an Nvidia T4 required to produce attribution maps. We showcase results on a ResNet50 network (He et al., 2016a) in Table 3, and provide additional results for ConvNext (Liu et al., 2022) and ViT (Dosovitskiy et al., 2020) in Table 5. All metrics are computed on a subset of the ImageNet validation set, made of $1280$ images. Importantly, this subset does not include the images used to compute $\sigma^{\star}$. Three key observations emerge from Table 3. First, FORGrad consistently boosts the faithfulness metrics for white-box methods (the quantitative improvement is validated by a statistical effect given by a Bayesian Anova analysis – section A.6). This observation experimentally validates the FORGrad method. Second, the $\mu$Fidelity metric shows little variations for most of the FORGrad white-box methods. We attribute this observation to the fact that the $\mu$Fidelity metric is strongly biased toward high-frequency content: random attribution maps with higher-frequency contents exhibit higher $\mu$Fidelity (see Table 4 and Fig. A.11). Third, in terms of computations, white-box (and gray-box) methods outperform black-box approaches. This increased efficiency stems from white-box attribution methods using more information per network-pass, thereby reducing the need for multiple network queries. Finally, we ensure that FORGrad does not affect the validity of white-box methods according to Adebayo et al. (2018) sanity check. We show in Fig. A.12 the resulting maps after randomizing the weights of a ResNet50 on white-box methods combined with FORGrad. The methods already failing the test before FORGrad still do, while the successful ones continue to pass the sanity check. All in all, FORGrad leverages the hidden information contained in the heatmaps provided by the white-box methods by filtering out the noisy high-frequencies in the gradient, making such methods rival with low-frequency black-box methods, as summarized in Fig. A.13.
@@
Table 2: Global ranking before (original) and after FORGrad. For each model, we show the five attribution methods with the highest metrics, before and after applying FORGrad. The explanation maps were computed on $1280$ images from the validation set of ImageNet, based on an aggregation of the three metrics computed by $F(\bm{x},\bm{\varphi})+\mu F(\bm{x},\bm{\varphi})$.
In Table 2, we aim to find the best attribution methods for ResNet50. To do so, we synthesize the faithfulness and $\mu$fidelity metrics into one single, comprehensive score. This score, labeled $F+\mu F$, corresponds to the sum of the faithfulness and the $\mu$fidelity. Table 2 reveals a notable shift in method rankings when integrating FORGrad. This change results in at least $4$ white-box methods in the top-5 position for the ResNet50 network. Similar findings for ConvNext and ViT are detailed in Table 6. The only black-box attribution method in the ranking is now the HSIC method. Note that HSIC demands 10 times more resources compared to its white-box counterparts.
@@ 6 Conclusion and Perspectives
This research started with the observation that white-box and black-box attribution methods produce attribution maps with different power spectra. Specifically, white-box methods show greater high-frequency power. We found that this high-frequency content originates from the gradient information used by white-box methods. We demonstrate that selectively filtering out these high frequencies does not alter the approximation error of the original gradient (using a first-order approximation). It suggests that such high-frequency content is non-informative and can be considered as noise. Further analysis revealed a strong connection between high frequencies in gradients and certain architectural constraints, including max-pooling operations. This led us to hypothesize that filtering out these high frequencies would enhance white-box attribution methods that rely on gradient information. To address this, we developed FORGrad, a low-pass filter with an optimal cut-off frequency tailored to each model architecture and attribution method. The application of FORGrad resulted in substantial improvements in XAI performance across all white-box attribution methods, challenging the traditional hierarchy of white-box and black-box methods. Notably, with FORGrad, white-box methods now dominate the top-5 rankings while also being more computationally efficient than their black-box counterparts.
@@
Our finding that the max-pooling operations significantly amplify high frequency in the gradient w.r.t input is surprising. This observation opens the doors to future research focused on designing novel pooling techniques that reduce high frequencies in the gradient itself. Additionally, conducting a systematic analysis of the Fourier spectrum of other architectural components, such as the normalization or attention layers, could lead to further improvements. Ideally, such an exploration would result in a comprehensive list of the contribution of each architectural element in the decision process. Studying their respective impact would contribute to building neural networks that are interpretable by design.
@@
So far, we have focused on gradient filtering to improve white-box attribution methods. One promising avenue would be to filter out high-frequencies in the gradient during training to reduce the presence of noisy artifacts during the optimization process and inherently improve its quality. We hypothesize that this approach may lead to better performance and faster convergence, as well as more qualitative information to explain the model using white-box methods. Networks equipped with such a filtering mechanism may also demonstrate enhanced adversarial robustness. We hope our work will spur interest in studying the frequency profile of networks, as it could lead to major improvement in terms of interpretability and robustness. Overall, we anticipate that our research will foster broader adoption of simpler and more efficient white-box methods for explainability, offering a better balance between faithfulness and computational efficiency.
@@ Acknowledgments
This work was supported to TS by ONR (N00014-19-1-2029), NSF (IIS-1912280 and EAR-1925481), DARPA (D19AC00015), NIH/NINDS (R21 NS 112743), the ANR-3IA Artificial and Natural Intelligence Toulouse Institute (ANR-19-PI3A-0004) to TS and RV. Additional support provided by the Carney Institute for Brain Science and the Center for Computation and Visualization (CCV). We acknowledge the Cloud TPU hardware resources that Google made available via the TensorFlow Research Cloud (TFRC) program as well as computing hardware supported by NIH Office of the Director grant S10OD025181.
Impact Statements
@@ A.1 Additional results
Rebuttal

### v4

@@ 5 FORGrad: FOurier Reparation of the Gradients
With FORGrad  we propose to filter out high-frequency content that is considered noise to improve the XAI performance of white-box attribution methods. We therefore apply a low-pass filter on the Fourier spectrum of the gradient of the classifier w.r.t the input (i.e. $\nabla_{\bm{x}}\bm{f}(\bm{x})$). This low-pass filter is equipped with a mechanism to adjust the cut-off frequency, $\sigma^{\star}$, based on the specific network architecture and attribution method.
@@
Fig. 5B illustrates the sensitivity of the faithfulness to changes in the cut-off frequency, $\sigma$, underscoring the necessity of selecting an optimal $\sigma^{\star}$. In Fig. 5A, we provide visual examples using the Gradient Input method (Ancona et al., 2018), where the gradient is low-pass filtered at various cut-off frequencies. One can observe that the initial noisy patterns (Fig. 5A, $\sigma=220$ ) progressively evolve into clearer and larger patches (Fig. 5A, $\sigma=80$), culminating in saliency maps that effectively emphasize the critical features for object recognition (Fig. 5A, $\sigma=10$). However, at very low $\sigma$ values (Fig. 5A, $\sigma=5$), the attribution map becomes overly diffuse, reducing interpretability. Fig. 5D compares the optimal $\sigma^{\star}$ values across different ResNetV2 (He et al., 2016b) architectures and attribution methods. We observe that even with the same attribution method, distinct architectures achieve their maximum faithfulness at different $\sigma^{\star}$ values. Likewise, the same network architecture can yield different optimal $\sigma^{\star}$ values when leveraging different attribution methods. This variability highlights the crucial role of a specific cut-off frequency, emphasizing that a one-size-fits-all approach may not be effective in maximizing the faithfulness of attribution methods in different contexts. Additionally, Fig. 5C presents a side-by-side comparison of attribution maps generated using the SmoothGrad method (Smilkov et al., 2017) (top-row) versus those refined with the FORGrad method (bottom row). A qualitative analysis shows that the FORGrad method yields more interpretable explanations, with a sharper focus on crucial features for object recognition. We refer the reader to Fig. A.12 for more examples of all white-box methods.
@@
White-box Saliency(Simonyan et al., 2013) 0.18 0.40 0.67 0.78
Saliency 0.28 0.39 0.53 0.89
Guidedbackprop(Ancona et al., 2018) 0.31 0.45 0.28 8.25
@@
Black & Gray-box GradCAM(Selvaraju et al., 2017a) 0.31 0.40 0.31 5.24
GradCAM++(Chattopadhay et al., 2018) 0.33 0.43 0.34 4.61
Occlusion(Ancona et al., 2018) 0.20 0.39 0.6 368
HSIC(Novello et al., 2022) 0.33 0.47 0.45 456
Sobol(Fel et al., 2021b) 0.34 0.47 0.47 578
RISE(Petsiuk et al., 2018) 0.41 0.34 0.55 626
Table 1: Results on Faithfulness and Stability metrics. Faithfulness, Fidelity, and Stability scores were obtained on 1,280 ImageNet validation set images, on an Nvidia T4 (For Faithfulness and Fidelity, higher is better). Time in seconds corresponds to the generation of 100 (ImageNet) explanations on an Nvidia T4. The first and second best results are in bold and underlined.
@@
FORGrad allows white-box methods to rival with black-box methods: We now evaluate the FORGrad method using XAI benchmarks, focusing on three established metrics: Faithfulness, $\mu$Fidelity, and Sensitivity (detailed in section  3). Our findings, presented in Table 1, compare the original performance of white-box attribution methods, their enhanced versions via the FORGrad method (marked with a $\star$), black-box, and gray-box methods. This comparison also includes the computational time measured on 100 images with an Nvidia T4 required to produce attribution maps. We showcase results on a ResNet50 network (He et al., 2016a) in Table 1, and provide additional results for ConvNext (Liu et al., 2022) and ViT (Dosovitskiy et al., 2020) in Table 4. All metrics are computed on a subset of the ImageNet validation set made of $1280$ images. Importantly, this subset does not include the images used to compute $\sigma^{\star}$. Three key observations emerge from Table 1. First, FORGrad consistently boosts the faithfulness metrics for white-box methods (the quantitative improvement is validated by a statistical effect given by a Bayesian Anova analysis – section A.6). This observation experimentally validates the FORGrad method. Second, the $\mu$Fidelity metric shows little variations for most of the FORGrad white-box methods. We attribute this observation to the fact that the $\mu$Fidelity metric is strongly biased toward high-frequency content: random attribution maps with higher-frequency contents exhibit higher $\mu$Fidelity (see Table 3 and Fig. A.13). However, the sensitivity metric, similarly to the faithfulness, is systematically improved for most models and methods, despite being evaluated with a $\sigma^{\star}$ optimized on the faithfulness. Third, in terms of computations, white-box (and gray-box) methods outperform black-box approaches. This increased efficiency stems from white-box attribution methods using more information per network pass, thereby reducing the need for multiple network queries. Additionally, we validate the optimality of our method by computing FORGrad on the attribution maps instead of the gradients. In Table 7, we compare the non-filtered metrics with the ones obtained with FORGrad on the attribution maps and FORGrad on the gradients (before computing the explanation). We observe a systematic improvement of the scores with FORGrad but more significant when applied to the gradients. Finally, we ensure that FORGrad does not affect the validity of white-box methods according to Adebayo et al. (2018) sanity check. We show in Fig. A.14 the resulting maps after randomizing the weights of a ResNet50 on white-box methods combined with FORGrad. The methods already failing the test before FORGrad still do, while the successful ones continue to pass the sanity check. All in all, FORGrad discards high-frequency artifacts to enhance low-frequency content of the attribution maps provided by the white-box methods by filtering out the noisy high-frequencies in the gradient, making such methods rival with low-frequency black-box methods, as summarized in Fig. A.15.
@@
[…]
5 GradCAM++ (0.42) GradCAM (0.42)
Table 2: Global ranking before (original) and after FORGrad. For each model, we show the five attribution methods with the highest metrics, before and after applying FORGrad. The explanation maps were computed on $1280$ images from the validation set of ImageNet, based on an aggregation of the three metrics computed by $(F(\bm{x},\bm{\varphi})+\mu F(\bm{x},\bm{\varphi})-S(\bm{x},\bm{\varphi})$.
In Table 2, we aim to find the best attribution methods for ResNet50. To do so, we synthesize the faithfulness, $\mu$fidelity, and sensitivity metrics into one single, comprehensive score. This score, labeled $(F+\mu F)-S$, corresponds to the sum of the faithfulness and the $\mu$fidelity, minus the sensitivity. Table 2 reveals a notable shift in method rankings when integrating FORGrad. This change results in at least $4$ white-box methods in the top-5 position for the ResNet50 network. Similar findings for ConvNext and ViT are detailed in Table 5.
@@ 6 Conclusion and Perspectives
This research started with the observation that white-box and black-box attribution methods produce attribution maps with different power spectra. Specifically, white-box methods show greater high-frequency power. We found that this high-frequency content originates from the gradient information used by white-box methods. We demonstrate that selectively filtering out these high frequencies does not alter the approximation error of the original gradient (using a first-order approximation). It suggests that such high-frequency content is non-informative and can be considered as noise. Further analysis revealed a strong connection between high frequencies in gradients and certain architectural constraints, including max-pooling operations. This led us to hypothesize that filtering out these high frequencies would enhance white-box attribution methods that rely on gradient information. To address this, we developed FORGrad  a low-pass filter with an optimal cut-off frequency tailored to each model architecture and attribution method. The application of FORGrad resulted in substantial improvements in XAI performance across all white-box attribution methods, challenging the traditional hierarchy of white-box and black-box methods. Notably, with FORGrad  white-box methods now dominate the top-5 rankings while also being more computationally efficient than their black-box counterparts.
Our current results demonstrate the efficiency of FORGrad on gradients containing artifacts. For this reason, the more high frequency is contained in the gradient, the more striking the improvement with FORGrad will be. Conversely, if filtering does not ameliorate the scores, FORGrad will not filter out anything, resulting in the original unfiltered explanation. Such a case is likely to happen with networks containing smooth gradients by design, such as 1-Lipschitz networks (Fel et al., 2022).
@@
Our finding that the max-pooling operations significantly amplify high frequency in the gradient w.r.t input is surprising. This observation opens the doors to future research focused on designing novel pooling techniques that reduce high frequencies in the gradient itself. Consistent with these results, the impact of FORGrad on ViT is less striking than on the Convolution model. This implies that a different phenomenon could happen in Transformer architectures and an investigation could lead to a better interpretability of this family of networks. Additionally, conducting a systematic analysis of the Fourier spectrum of other architectural components, such as the normalization or attention layers, could lead to further improvements. Ideally, such an exploration would result in a comprehensive list of each architectural element’s contribution to the decision process. Studying their respective impacts would contribute to building neural networks that are interpretable by design.
The impact of FORGrad is measured by the standard metrics of XAI. We evaluated FORGrad on three different metrics and observed a clear improvement in two of them. However, finding appropriate metrics is still an open question in this field, and these results could be affected by the development of new metrics. In Bhatt et al. (2020), the authors advocate for a way to combine several XAI metrics to obtain the best method. They propose an aggregate of Faithfulness, Sensitivity, and Complexity of the explanation, arguing that the best explanation should maximize the first and minimize the latter two. Here, we evaluate FORGrad on two of these metrics and complete this framework with the Computational cost measured by the time in seconds. We, therefore, believe that FORGrad remains an optimal solution to compute faithful explanations in a reasonable amount of time, given the current metrics to evaluate the quality of an explanation. More interestingly, one possible explanation for why black-box methods tend to outperform white-box methods is that their optimization objectives align with the evaluation metrics (e.g., Occlusion with Deletion metric). In contrast, the gradients result from optimizing the model for a classification task. Therefore, it is encouraging to achieve scores comparable to those of black-box methods using white-box attribution methods with FORGrad  bringing Saliency and other gradient-based methods back into prominence.
@@
So far, we have focused on gradient filtering to improve white-box attribution methods. One promising avenue would be to filter out high-frequencies in the gradient during training to reduce the presence of noisy artifacts during the optimization process and inherently improve its quality. We hypothesize that this approach may lead to better performance and faster convergence, as well as more qualitative information to explain the model using white-box methods. Networks equipped with such a filtering mechanism may also demonstrate enhanced adversarial robustness. We hope our work will spur interest in studying network frequency profiles, as it could lead to major improvements in terms of interpretability and robustness. Overall, we anticipate that our research will foster broader adoption of simpler and more efficient white-box methods for explainability, offering a better balance between faithfulness and computational efficiency.
@@ Acknowledgments
This work was supported by grants from ONR (N00014-19-1-2029) and NSF (IIS-1912280 and EAR-1925481) to TS and RV and the ANR-3IA Artificial and Natural Intelligence Toulouse Institute (ANR-19-PI3A-0004). The Carney Institute for Brain Science and the Center for Computation and Visualization (CCV) provided additional support. We acknowledge the Cloud TPU hardware resources that Google made available via the TensorFlow Research Cloud (TFRC) program and computing hardware supported by NIH Office of the Director grant S10OD025181.
Impact Statement
@@ A.1 Additional results
Frequency power per category.
Figure A.1: 2D Fourier spectrum of attribution methods. We show on the top row the Fourier spectrum of white-box attribution methods and of the gray and black-box methods on the bottom row, computed with a ResNet50V2 (each spectrum is normalized to highlight values ranging within the same boundaries). The three families can be distinguished by methods but also by their signature in the Fourier domain. The former method has magnitudes largely concentrated in the low frequencies, while the latter is more spread out: it features non-trivial magnitudes almost everywhere, including in high frequencies. (Blue = low power, Green = high power, Red = very high power)
A.2 Gradient approximation with ConvNeXt and ViT
In this section, we show the results of the gradient approximation presented in the main text on ConvNeXt and ViT. For each network, we evaluate the importance of high-frequency content in the gradient using a first-order approximation of the model, i.e. $\bm{f}(\bm{x}+\bm{\varepsilon})\approx\bm{f}(\bm{x})+\bm{\varepsilon}\nabla_{\bm{x}}\bm{f}(\bm{x})$, and we compute the $\ell_{2}$ approximation error, i.e. $\lVert\bm{f}(\bm{x}+\bm{\varepsilon})-(\bm{f}(\bm{x})-\epsilon.\nabla_{\sigma})\rVert_{2}/\lVert\bm{f}(\bm{x}+\bm{\varepsilon})-(\bm{f}(\bm{x})-\epsilon.\nabla_{\sigma_{max}})\rVert_{2}$. We additionally provide four control conditions representing the absence of relevant information in the gradient using a gradient containing permuted content, zeros, uniform noise, or random 2D gaussians. The left plot shows the results on ConvNeXt and the right plot for ViT. The reader can notice that compared to Fig. 4.A, the approximation errors of the permuted and zero control conditions reach an order of magnitude more than the most filtered gradient, confirming our precedent results with ResNet50. The 2D Gaussian control condition represents information of low-frequency content under the form of randomly located 2D Gaussians instead of the gradient. Contrary to the “zero” condition (also representing low-frequency information), the activity is not null and might be more informative. However, it reaches the same approximation as the other baseline, being at least twice as high as the more filtered gradient. These additional results confirm the hypothesis that high frequencies in the gradient can be considered not as informative as the low frequencies.
Figure A.2: Gradient approximation experiment on ResNet50, ConvNeXt and ViT. We evaluate the importance of high-frequency content in the gradient using a first-order approximation of the model, i.e. $\bm{f}(\bm{x}+\bm{\varepsilon})\approx\bm{f}(\bm{x})+\bm{\varepsilon}\nabla_{\bm{x}}\bm{f}(\bm{x})$, and we compute the $\ell_{2}$ approximation error, i.e. $\lVert\bm{f}(\bm{x}+\bm{\varepsilon})-(\bm{f}(\bm{x})-\epsilon.\nabla_{\sigma})\rVert_{2}/\lVert\bm{f}(\bm{x}+\bm{\varepsilon})-(\bm{f}(\bm{x})-\epsilon.\nabla_{\sigma_{max}})\rVert_{2}$. Gradients in which we remove high-frequency content (dark blue) produce an error closer to the baseline (green), compared to the control conditions (purple).
Downsampling operations in a ResNet50V2.
Downsampling operations in a VGG16.
Figure A.8: Visualization of the gradient (w.r.t. the input) probed at different layers. We observe that both in the VGG MaxPooling and the VGG untrained, a checkerboard signal emerges in the gradients. Such a signal might cause an increased power in high frequencies. In contrast, there is no such signal in the VGG Average Pooling. It suggests that MaxPooling operations, creating such a checkerboard pattern, might contaminate the gradients with high-frequency artifacts.
Finally, we vary the stride value of the MaxPooling operations in a trained VGG16 with MaxPooling2D. In Fig A.9, we find that performing a MaxPooling with no stride (meaning using a value of 1) reduces the amount of high-frequency content introduced per layer, compared to using a stride value superior to 1. However, increasing the stride value from 2 to 4 does not change significantly the amount of high-frequency content.
Figure A.9: Effect of the stride value of MaxPooling in the VGG16. We observe that performing a MaxPooling with a stride of 1 (akin to no stride), reduces the amount of high-frequency artifacts included in the gradients of the model. Varying the stride value (2 or 4) does not change the high-frequency power.
A.4 The importance of an adaptive sigma
In Fig. A.10, we show the evolution of the faithfulness w.r.t the value of $\sigma$. We select the optimal cutoff frequency $\sigma\star$ by maximizing the faithfulness. These values range from $224$ to $2$ (we plot the values here starting at $160$ for the sake of visualization). We can observe that in most cases, the faithfulness score increases significantly after filtering the gradient. We also observe a pattern of evolution similar for the methods belonging to the same sub-categories (i.e. SmoothGrad, SquareGrad, VarGrad). Additionally, the filtering is more useful for convolutional networks (ResNet50V2, ConvNeXt) than for ViT. This can be explained by the absence of MaxPooling or striding in transformer architectures and smoother gradients. In Fig. A.11, we plot the fidelity metric using the same $\sigma$ value as in Fig. A.10. The pattern of evolution is significantly different, and the $\sigma$ value maximizing the faithfulness metric does not always correspond to the optimal $\sigma$ value for the fidelity metric. However, it is possible to optimize $\sigma$ on the fidelity metric as there is at least one value per model and method which is increased compared to the absence of filtering.
Figure A.11: Evolution of the fidelity depending on the cut-off frequency $\sigma$. The faithfulness evolves differently per methods and models depending on the cut-off frequency $\sigma$. Consistently, filtering improves the faithfulness score or at least does not reduce it.
A.5 Qualitative examples
A.6 Control experiments
We provide in this section some control experiments on FORGrad.
Controlling the metrics.
Method name Deletion Insertion Faithfulness $\mu$ Fidelity
Saliency 0.127 0.316 0.189 0.390
Random low-freq. 0.205 0.187 -0.018 0.140
Random high-freq. 0.138 0.179 0.041 0.310
Sanity check for Saliency.
Statistical effect of FORGrad.
To show the statistical effect of FORGrad  for each metric we performed a Bayesian ANOVA (see details below). We analyzed the metrics by means of Bayesian ANOVAs, considering BEFORE-AFTER and ATTRIBUTION-METHODS as fixed factors, and MODELS as a random term. In all analyses, we computed Bayes Factors (BF) as the ratio between the models testing the alternative against the null hypothesis. The alternative hypothesis states that the factor is important for explaining the variance in the data (i.e., there is a significant difference between conditions, for example before and after). All BFs are denoted as BF10. In practice, BFs provide substantial (BF¿ 5) or strong (BF¿ 10) evidence in favor of the alternative hypothesis, and low BF (BF¡ 0.5) suggests a lack of effect (Masson, 2011). Confirming our results, we found a strong difference in the BEFORE-AFTER factor (comparing scores before and after FORGrad) in the INS metric (BEFORE-AFTER factor BF10=11.33, error= 1%), and moderate evidence for the DEL metric (BEFORE-AFTER factor BF10=1.79, error= 2%). All in all, considering the small number of samples used for this analysis, we are very confident that the results of the statistical analysis confirm a significant difference between the scores obtained before and after applying FORGrad on the different attribution methods.
Standard Deviation.
We additionally include in Table 7 the standard deviation computed on the scores obtained on each image, for each attribution and metric on ResNet50.
A.7 Quantitative results of FORGrad
Illustrating the contribution of FORGrad.
Figure A.15: Effect of FORGrad on white-box attribution methods. Before applying FORGrad  most white-box methods (reddish points) contain a lot of power in the high-frequencies (x-axis) but a lower explainability score (y-axis) than other types of methods (blue and green points). FORGrad corrects white-box methods by low-pass filtering the gradient (reducing the high-frequency power), bringing them back to a faithfulness closer to gray and black-box methods.
We additionally report here the faithfulness and fidelity scores of white-box methods before and after FORGrad  and compare them with gray and black-box methods for ResNet50V2, ConvNeXt, and ViT. GradCAM methods can’t be tested on ViT because they are based on convolution so are limited to CNNs. For the $3$ models, before FORGrad  the best methods are always black-box. Even though FORGrad does not allow to systematically beat them, it brings white-box methods back to competitive performance. Considering their computational efficiency, FORGrad rends white-box methods very attractive given the accuracy/efficiency tradeoff.
ResNet50 ConvNeXT ViT
Faith.($\uparrow$) $\mu$Fid.($\uparrow$) Stab.($\downarrow$) Time($\downarrow$) Faith.($\uparrow$) $\mu$Fid.($\uparrow$) Stab.($\downarrow$) Time($\downarrow$) Faith.($\uparrow$) $\mu$Fid.($\uparrow$) Stab.($\downarrow$) Time($\downarrow$)
Gradient-based Saliency(Simonyan et al., 2013) 0.18 0.40 0.67 0.78 -0.01 0.46 7e-05 1.89 0.09 0.41 0.06 0.45
Saliency 0.27 0.39 0.53 0.89 0.06 0.47 3e-05 1.64 0.10 0.40 0.03 0.46
Guidedbackprop(Ancona et al., 2018) 0.31 0.45 0.28 8.25 -0.02 0.45 4e-05 19.3 0.05 0.39 0.04 5.48
[…]
GradInput(Ancona et al., 2018) 0.20 0.36 0.42 0.73 0.05 0.43 0.005 1.93 0.16 0.40 0.02 0.53
[…]
Int.Grad(Sundararajan et al., 2017) 0.24 0.39 0.72 42.7 0.11 0.49 0.16 93.5 0.19 0.41 0.12 24.6
[…]
SmoothGrad(Smilkov et al., 2017) 0.23 0.45 0.22 46.6 0.02 0.45 4e-05 104 0.11 0.45 0.03 26.4
[…]
VarGrad(Seo et al., 2018) 0.36 0.46 0.003 41.5 -0.09 0.21 1e-10 93.7 0.16 0.39 0.001 27.1
[…]
SquareGrad(Seo et al., 2018) 0.36 0.45 0.003 42.1 -0.02 0.33 2e-09 93.4 0.17 0.41 0.001 20.5
[…]
Prediction-based GradCAM(Selvaraju et al., 2017a) 0.31 0.40 0.31 5.24 0.03 0.26 0.005 6.80 n.a n.a n.a n.a
GradCAM++(Chattopadhay et al., 2018) 0.33 0.43 0.34 4.61 0.02 0.19 0.001 7.06 n.a n.a n.a n.a
Occlusion(Ancona et al., 2018) 0.17 0.39 0.60 368 0.05 0.45 0.05 782 0.09 0.37 0.09 183
HSIC(Novello et al., 2022) 0.33 0.47 0.45 456 0.08 0.54 0.14 1358 0.20 0.47 0.14 248
Sobol(Fel et al., 2021b) 0.34 0.47 0.47 578 0.12 0.53 0.16 1156 0.14 0.47 0.05 342
RISE(Petsiuk et al., 2018) 0.41 0.34 0.55 626 0.13 0.49 0.15 1534 0.18 0.44 0.08 487
ResNet50 ConvNeXT ViT
Original FORGrad Original FORGrad Original FORGrad
[…]
Table 5: Global ranking before (original) and after FORGrad. For each model, we show the 5 attribution methods with the highest metrics, before and after applying FORGrad. The explanation maps were computed on $1280$ images from the validation set of ImageNet, based on an aggregation of the three metrics computed by $F(\bm{x},\bm{\varphi})+\mu F(\bm{x},\bm{\varphi})-S(\bm{x},\bm{\varphi})$.
Additional quantitative control measurements.
We finally include additional experiments to control the robustness of FORGrad on the different metrics and parameters. Table 6 summarizes the results of FORGrad when the baseline value of Insertion and Deletion metrics is changed from 0 to uniform noise between -1 and 1. We can see that the scores indeed differ, but the effectiveness of the method does not. Table 7 evaluates the effect of applying FORGrad (i.e. low-pass filtering using an optimal cut-off value) after computing the attribution map instead of filtering the gradient prior to generating the attribution map ($\dagger$ symbol) on ResNet50. Even though the metrics improve compared to the absence of filtering, filtering the gradient before computing the attribution map is the most optimal.
ResNet50
Del.($\downarrow$) Ins.($\uparrow$) Faith.($\uparrow$)
White-box Saliency(Simonyan et al., 2013) 0.08 0.27 0.17
Saliency 0.09 0.37 0.28
Guidedbackprop(Ancona et al., 2018) 0.05 0.33 0.28
[…]
GradInput(Shrikumar et al., 2017) 0.08 0.23 0.15
[…]
Int.Grad(Sundararajan et al., 2017) 0.11 0.17 0.06
[…]
SmoothGrad(Smilkov et al., 2017) 0.13 0.31 0.18
[…]
VarGrad (Adebayo et al., 2018) 0.12 0.28 0.16
[…]
SquareGrad(Seo et al., 2018) 0.12 0.29 0.17
[…]
Black & Gray-box GradCAM(Selvaraju et al., 2017a) 0.10 0.40 0.30
GradCAM++(Chattopadhay et al., 2018) 0.08 0.41 0.33
Occlusion(Ancona et al., 2018) 0.08 0.28 0.20
HSIC(Novello et al., 2022) 0.06 0.42 0.36
Sobol(Fel et al., 2021b) 0.07 0.40 0.33
RISE(Petsiuk et al., 2018) 0.07 0.41 0.34
Table 6: Results on Deletion, Insertion, and Faithfulness metrics. Deletion, Insertion, and Faithfulness scores were obtained on 1,280 ImageNet validation set images, on an Nvidia T4 (For Insertion and Faithfulness, higher is better). The scores that are improved by FORGrad  are in bold